OpenStudy (anonymous):

Divide and simplify. y^8 / y^13=

OpenStudy (anonymous):

We can use the rule of exponennets : a^m/a^n=a^m-n

OpenStudy (anonymous):

1 / y^5

OpenStudy (anonymous):

1/y^5 y^-5

OpenStudy (anonymous):

thus y^8 / y^13=y^-5